Journalist Report – February 28th

Sol 7 – From Interstellar to gravitropism experiment

Author : Nicolas Wattelle

Rest has been great (we all appreciated Christopher Nolan’s movie), now it is time to go back to efficient work! And I think we did: our schedule was not really full so we took advantage of it to prepare the next days, and to do what we usually have not the time to. The upcoming days are very exciting, this is partly due to the weather improvement, so we can broaden our EVA plans and make the mission go forward.

During the day, our focus was thus on planning, data processing and maintenance. For the planning, discussions have taken place during the whole early afternoon on the upper deck, and one of the great news is concerning the geological field study (mainly for Marine and Léa who are in charge of it) that might begin no later than tomorrow. Cerise, our commander, make a great job everyday to handle the tricky moving planning, dealing with all constraints (even the last-minute ones), and she always tries to satisfy each of us’ interests. She is a great commander (honestly, without any will to become the favourite).

In the middle of all this, two remarkable novelties: Elena has started the preparation of a spirulina culture; and Mathéo tried for the first time to pilot the drone from inside the Hab during an EVA. Léa who was Habcom at this moment, found this very useful to be able to follow the astronauts this way, it is much easier to understand their progression with this point of view (can you spot the astronauts on the today’s drone shot?).

Also, maintenance has been performed on the Solar Musk cupola, on the field mill (we had a mishap with it during the last EVA) and on the gravitropism experiment.

But what is gravitropism? Plants grow, we already know that, but in fact, they grow in response to stimuli. The stimulus can be various things: light, water presence, heat, fertilizer… In these cases, we respectively speak about phototropism, hydrotropism, thermotropism and chemotropism. Each plant is sensitive to these phenomena, like in a competition between the latter, and each plant is more or less responsive to a phenomenon. We know since the XVIIIth century that plants react to gravity. Gravitropism (and so gravity) is one of the reasons why roots go down and rods go up.

During our mission, and in collaboration with middle and high school students, we try to observe and disturb gravitropism. How do we do this? Thanks to inertia! We grow plants (ones with rapid growth, like lentils, beans, wheat…), on a turning wheel (almost 30 cm of radius), in order to make seeds feel centrifugal force. This way, when you add the contribution of this force to gravitational force, you obtain an inclined resultant (the force is not purely vertical anymore). And we might be able to see modifications in the growth. This is the idea of an historic experiment named “Knight’s Wheel”.

We had some trouble doing this experiment during our first week, but today, we took this experiment back to zero, hoping it will show interesting things before next weekend. Plus, we plan to experiment another unperformed protocol during our last week: what if we turn the wheel vertical? We are very curious to produce data on it!

PS: In addition, you can find in the daily pictures the M81 galaxy 11,74 million of light years away, observed by Marine. A picture like this might leave us dreamers, imagining a solar system similar to ours, and who knows… An inhabited planet like ours?

EVA10 Report – February 28th

Crew 263 EVA Report 28-02-2022

EVA # 10

Author: Elena López-Contreras González

Purpose of EVA: Augmented Reality application testing with a HoloLens device and drone image capture of the station.

Start time: 10:30h

End time: 10:51h

Narrative: EVA Participants entered the airlock at 10:25h. Depressurization ended at 10:30h and the participants exited the airlock. Elena placed the drone on the ground and checked Perseverance.

The EVA started 10 minutes later than scheduled due to troubleshooting the application. Nevertheless, the re-enter time is well within the scheduled limit.

Perseverance: Usage Duration – 228.1 Battery Charge – 100%

The drone was correctly deployed from the inside of the Hab. Several videos and pictures were taken, most notably a Birdseye view of the station, a view with the antenna and a desert view.

Following the same route taken in previous EVAs, the first waypoint was successfully placed on the RAM. The pictures were successfully taken as well by Cerise Cuny, the astronaut wearing the HoloLens. The following waypoints were placed in the Hab, GreenHab, Science Dome and Solar Observatory. The pictures were successfully taken in every point.

Once the waypoints had been placed, the second tour of the station began. The visible waypoints were reported back to HabCom and will be analyzed.

At 10:43, after finishing the second tour of the station, the drone was retrieved –the pilot posed it on the ground for safe retrieval-.

The astronauts entered the airlock at 10:46h and pressurization began. At 10:51h they both entered the Hab.

Nominal EVA and the application was successfully tested.

Coordinates (use UTM NAD27 CONUS): 4250750 N, 518250 E

Participants: Elena López-Contreras and Cerise Cuny. Léa Rouverand as HabCom.

Road(s) and routes per MDRS Map: NaN

Mode of travel: Foot

Vehicles used (If applicable): NaN

EVA9 Report – February 28th

Crew 263 EVA Report 28-02-2022

EVA # 9

Author: Valentine Bourgeois

Purpose of EVA: Atmospheric Experiment Battery Change

The atmospheric experiment includes 5 devices: LOAC (Light Optical Aerosol Counter), an Electrical Field Mill, Purple Air (Aerosol Counter), Mega-Ares (Electrical Field Counter) and a Weather station. The aim of this experiment is to do testing of Martian-Atmospheric Study Equipment and their maintenance by human operators. Performing their maintenance and set-up while wearing spacesuits and following EVA protocols is the purpose of the experiment. All devices are located at the same place.

Batteries for the LOAC, the Electrical Field Mill, Purple Air and Mega-Ares needs to be changed daily. The Electrical Field Mill and Mega-Ares share the same battery. Hence, during this maintenance EVA, the 3 batteries needed to be replaced.

Start time: 08:58

End time: 09:35

Narrative: EVA Participants entered the airlock at 08:58. Depressurization ended at 09:03. EVA participants checked the rover batteries just in case. The obtained information is:

Perseverance: Usage Duration – 228.1 Battery Charge – 100%

Spirit: Usage Duration – 157.2 Battery Charge – 100%

Opportunity: Usage Duration – 84.7 Battery Charge – 100%

Curiosity: Usage Duration – 168.7 Battery Charge – 100%

The EVA participants arrived at the Atmospheric Experiment location at 09:09. The 3 batteries were replaced. Lea and Valentine changed the LOAC battery and changed the SD card to collect data. Valentine changed the Purple Air Battery and Nicolas changed the Electrical Field Mill and Mega-Ares batteries. The status of the devices was checked and we realized that the Electrical Field Mill was not working anymore. The three of us have dismounted it to bring it back to the station for maintenance.

At 09:25, the participants left the atmospheric experiment location.

At 09:29, the participants arrived at the Hab and entered Airlock.

09:35, pressurization over.

Destination: Atmospheric experience site (at intersection between Cow Dung Road and Entrance Road)

Coordinates (use UTM NAD27 CONUS): 518288 E, 4250586 N

Participants : Valentine Bourgeois (EVA leader), Nicolas Wattelle, Léa Rouverand. HabCom was Marine Prunier.

Road(s) and routes per MDRS Map: Entrance Road

Mode of travel: Foot

EVA Report – February 28th

Crew 263 EVA Report 28-02-2022

EVA # 9

Author: Valentine Bourgeois

Purpose of EVA: Atmospheric Experiment Battery Change

The atmospheric experiment includes 5 devices: LOAC (Light Optical Aerosol Counter), an Electrical Field Mill, Purple Air (Aerosol Counter), Mega-Ares (Electrical Field Counter) and a Weather station. The aim of this experiment is to do testing of Martian-Atmospheric Study Equipment and their maintenance by human operators. Performing their maintenance and set-up while wearing spacesuits and following EVA protocols is the purpose of the experiment. All devices are located at the same place.

Batteries for the LOAC, the Electrical Field Mill, Purple Air and Mega-Ares needs to be changed daily. The Electrical Field Mill and Mega-Ares share the same battery. Hence, during this maintenance EVA, the 3 batteries needed to be replaced.

Start time: 08:58

End time: 09:35

Narrative: EVA Participants entered the airlock at 08:58. Depressurization ended at 09:03. EVA participants checked the rover batteries just in case. The obtained information is:

Perseverance: Usage Duration – 228.1 Battery Charge – 100%

Spirit: Usage Duration – 157.2 Battery Charge – 100%

Opportunity: Usage Duration – 84.7 Battery Charge – 100%

Curiosity: Usage Duration – 168.7 Battery Charge – 100%

The EVA participants arrived at the Atmospheric Experiment location at 09:09. The 3 batteries were replaced. Lea and Valentine changed the LOAC battery and changed the SD card to collect data. Valentine changed the Purple Air Battery and Nicolas changed the Electrical Field Mill and Mega-Ares batteries. The status of the devices was checked and we realized that the Electrical Field Mill was not working anymore. The three of us have dismounted it to bring it back to the station for maintenance.

At 09:25, the participants left the atmospheric experiment location.

At 09:29, the participants arrived at the Hab and entered Airlock.

09:35, pressurization over.

Destination: Atmospheric experience site (at intersection between Cow Dung Road and Entrance Road)

Coordinates (use UTM NAD27 CONUS): 518288 E, 4250586 N

Participants : Valentine Bourgeois (EVA leader), Nicolas Wattelle, Léa Rouverand. HabCom was Marine Prunier.

Road(s) and routes per MDRS Map: Entrance Road

Mode of travel: Foot

Operations Report – February 28th

Crew 263 Operations Report 28-02-2022

SOL: 7

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA

ROVERS

Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Opportunity rover used: No

Hours: 84.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Curiosity rover used: No

Hours: 68.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Perseverance rover used: No

Hours: 228.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

General notes and comments: NA

Summary of Hab operations:

WATER USE: 20 gallons

Water (static tank): 397 gallons

Water (loft tank): 35 gallons

Water Meter: 01561697 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No

Summary of internet: NA

Summary of suits and radios: I tested the malfunctioning earpiece with every radio and through different modes but it is not working anymore

Summary of GreenHab operations: NA

Summary of ScienceDome operations:

Dual split: Off

SOC: 87% at 5:45pm which seems normal thanks to Shannon’s intervention on it

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO Report

Questions, concerns and requests to Mission Support: NA

Sol Summary – February 28th

Crew 263 Sol Summary Report 28-02-2022

Sol: 7

Summary Title: Back to work

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: Surprisingly after a day of rest, we felt more tired this morning when we woke up than during the previous week. The two morning EVAs went smoothly. Meanwhile, the crew Engineer and I tested the broken earpiece with several radios and modes. We spotted some exposed wires which may be the source of the issue. As the earpiece has not been subject to any violent/unusual use, maybe it is very sensitive. To make sure we do not damage any other earpiece, we will think about ways to attach the wire so that it cannot get caught in anything while on EVA. Our astronomer worked on astrophotography images and tried to retrace the light curve of an exoplanet transit using the data from the New Mexico remote telescope.

We had lunch all together, as usual: we deeply value these moments when we can share the latest development in our experiments, our feelings about being on Mars, far from our family and friends, how confinement and isolation impacts our mood, etc.

In the afternoon, our GreenHab officer started a spiruline experiment, which reminded her of her role even with the GreenHab being closed. We prepared our EVA plans for tomorrow. Some of us did artistic activities (water color painting and knitting).

Anomalies in work: None.

Weather: Sunny, 12°C at 6pm (warmer than last week).

Crew Physical Status: Good.

EVA: 2 EVAs (n°8,9). Both went well, no issue to report.

Reports to be filed: Sol Summary, 2 EVA Requests, 2 EVA Reports, Operations Report, Journalist Report, HSO Report.

Support Requested: No support requested.